# Reasoning-Enhanced Object-Centric Learning for Videos

### v1

@@
Abstract
Object-centric learning aims to break down complex visual scenes into more manageable object representations, enhancing the understanding and reasoning abilities of machine learning systems toward the physical world. Recently, slot-based video models have demonstrated remarkable proficiency in segmenting and tracking objects, but they overlook the importance of the effective reasoning module. In the real world, reasoning and predictive abilities play a crucial role in human perception and object tracking; in particular, these abilities are closely related to human intuitive physics. Inspired by this, we designed a novel reasoning module called the Slot-based Time-Space Transformer with Memory buffer (STATM) to enhance the model’s perception ability in complex scenes. The memory buffer primarily serves as storage for slot information from upstream modules, the Slot-based Time-Space Transformer makes predictions through slot-based spatiotemporal attention computations and fusion. Our experiment results on various datasets show that STATM can significantly enhance object-centric learning capabilities of slot-based video models.
1 Introduction
@@
Figure 1: Slot-based Time-Space Transformer with Memory buffer architecture overview.
@@
The two main steps of such a model operate in a positive feedback loop. The more accurate the predictions, the better the corrections become. Consequently, the more accurate the corrections, the more precise the information provided for the prediction step is, leading to better predictions. Therefore, having a reasonable and efficient predictor is crucial for the entire model. In real-world scenarios, humans also engage in prediction as a crucial aspect of their object perception and tracking, but their prediction behaviors often involve more intricate processes. Humans typically combine the motion state of an object with the interactions of other objects to predict possible future states and positions of the object. The object’s motion state is inferred by humans using their common sense from the object’s past positions over a while. In so doing, humans enhance their ability to recognize and track relevant objects within complex scenes, which is an integral component of human intuitive physics (Sudderth, 2006; Ullman et al., 2017; Mitko & Fischer, 2020). The prediction step in SAVi and SAVi++ is similar to human inference, but their predictor module is somewhat simplistic, as it relies solely on single-frame information from the current time step for prediction.
@@
Drawing inspiration from human behavior, we introduce a novel prediction module aimed at enhancing slot-based models for video. This module comprises two key components: 1) Slot-based Memory Buffer: primarily designed to store historical slot information obtained from the upstream modules; and 2) Slot-based Time-Space Transformer Module: designed by applying spatiotemporal attention mechanisms to slots for inferring the temporal motion states of objects and calculating spatial objects interactions, which also integrates time and space attention results. We term the proposed model as Slot-based Time-Space Transformer with Memory buffer (STATM). Upon substituting the prediction module of SAVi and SAVi++ into the STATM, we observe distinct impacts of different spatiotemporal fusion methods on SAVi and SAVi++. By employing an appropriate fusion method and memory buffer sizes, we observed a significant enhancement in the object segmentation and tracking capabilities of SAVi and SAVi++ on videos containing complex backgrounds and multiple objects per scene.
2 Related Work
@@
Object-centric Learning. In recent years, object-centric learning has emerged as a significant research direction in computer vision and machine learning. It aims to enable machines to perceive and understand the environment from an object-centered perspective, thereby constructing more intelligent visual systems. There is a rich literature on this research, including SQAIR (Kosiorek et al., 2018), R-SQAIR (Stanić & Schmidhuber, 2019), SCALOR (Jiang et al., 2019), Monet (Burgess et al., 2019), OP3 (Veerapaneni et al., 2020), ViMON (Weis et al., 2020), PSGNet (Bear et al., 2020), SIMONe (Kabra et al., 2021), and others (Kahneman et al., 1992; Kipf et al., 2019; Zhang et al., 2022; Xie et al., 2022). Slot-based Models represent a prominent approach within object-centric learning. They achieve this by representing each object in a scene as an individual slot, which is used to store object features and attributes (Locatello et al., 2020; Kumar et al., 2020; Zoran et al., 2021; Singh et al., 2021; Yang et al., 2021; Zoran et al., 2021; Ye et al., 2021; Hassanin et al., 2022; Wang et al., 2023; Heravi et al., 2023; Wu et al., 2023).
Prediction and Inference on Physics. The implementation of object-centric physical reasoning is crucial for human intelligence and is also a key objective in artificial intelligence. Interaction Network (Battaglia et al., 2016) as the first general-purpose learnable physics engine, is capable of performing reasoning tasks centered around objects or relationships. Another similar study is the Neural Physics Engine (Chang et al., 2016). On the other hand, Visual Interaction Networks (Watters et al., 2017) can learn physical laws from videos to predict the future states of objects. Additionally, there are many models developed based on this foundation (Henderson & Lampert, 2020; Chen et al., 2021; Dittadi et al., 2021; Jusup et al., 2022; Meng et al., 2022; Piloto et al., 2022; Singh et al., 2022; Driess et al., 2023). In order to achieve a deeper understanding of commonsense intuitive physics within artificial intelligence systems, (Piloto et al., 2022) have built a system capable of learning various physical concepts, albeit requiring access to privileged information such as segmentation. Our research primarily aims to construct an object-centric system for object perception, learning of physics, and reasoning.
Figure 2: Left: Fusion approaches of spatiotemporal attention explored in our study. (a) The sum of computed temporal attention and spatial attention (T+S). (b) Spatial attention computation followed by using the outcome as input for temporal attention (ST). (c) Temporal attention computation followed by using the outcome as input for spatial attention (TS). Right: Spatiotemporal attention computation architectures explored in our study. The green slots represent those employed for spatial attention computation, while the orange slots are indicative of those used for temporal attention computation. (d) Corresponding Slot Attention (CS). (e) All Slot Attention (AS).
@@
Slot-based Attention and spatiotemporal Attention. Our current work is closely related to slot-based attention and spatiotemporal attention. There are a lot of works related to slot-based attention (Locatello et al., 2020; Hu et al., 2020; Kumar et al., 2020; Zoran et al., 2021; Singh et al., 2021; Yang et al., 2021; Zoran et al., 2021; Ye et al., 2021; Hassanin et al., 2022; Wang et al., 2023; Heravi et al., 2023; Wu et al., 2023). Spatiotemporal attention mechanisms are particularly effective in handling video data or time-series data, allowing networks to understand and leverage relationships between different time steps or spatial positions (Li et al., 2020; Bertasius et al., 2021; Luo et al., 2021). Currently, they find wide applications in various fields such as video object detection and tracking (Lin et al., 2021; Chen et al., 2022), action recognition (Yang et al., 2022), natural language processing (Xu et al., 2020; Weld et al., 2022), medical image processing (Zhang et al., 2020), among many others (Ding et al., 2020; Yuan et al., 2020; Cheng et al., 2020).
@@
3 Slot-based Time-Space Transformer with Memory Buffer (STATM)
@@
To enhance the slot-based video models, we introduce a new module called the Slot-based Time-Space Transformer with Memory Buffer (STATM) as the predictor. STATM is primarily designed to support causal reasoning and prediction for object-centric downstream tasks based on slots. This module consists of two key components: 1) the memory buffer, and 2) the Slot-based Time-Space Transformer (STAT). The memory buffer serves as a repository for storing historical slot information obtained from upstream modules, while STAT utilizes the information stored in the memory buffer for prediction and causal reasoning.
3.1 Memory Buffer
@@
3.2 Slot-based Time-Space Transformer
The primary role of STAT (Slot-based Time-Space Transformer) lies in leveraging slot data from the memory buffer to facilitate slot-based dynamic temporal reasoning and spatial interaction computations. Furthermore, it integrates the outcomes of temporal reasoning and spatial interactions to achieve a unified understanding. Specifically, for temporal dynamic reasoning, a cross-attention mechanism is employed, which effectively utilizes historical context stored in the memory buffer to enable accurate predictions of future states. Meanwhile, for spatial interaction computations, we employ a self-attention mechanism that operates on slot representations to compute the relevance between different slots within the $S$. The results obtained from temporal dynamic reasoning and spatial interaction computation are merged to provide a holistic understanding encompassing both temporal dynamics and spatial interactions. This comprehensive representation enhances the model’s capability for accurate prediction and reasoning in object-centric tasks. We propose three approaches as illustrated in  Figure 2 a-c.
We also introduced two computational architectures for spatiotemporal attention as illustrated in  Figure 2 d-e. (1) Corresponding Slot Attention (CS): For slot $s_{(i,t)}$, temporal attention is computed by using it and corresponding slots in $\{s_{(i,0)},\dots,s_{(i,t-1)}\}$, while spatial attention computation is performed using it and all slots within $\{s_{(0,t)},\dots,s_{(N,t)}\}$. (2) All Slot Attention (AS): For slot $s_{(i,t)}$, temporal attention is computed by using it and all slots in $\{s_{(0,0)},\dots,s_{(N,t-1)}\}$. The spatial attention computation remains the same as in approach CS.
@@ 3.2 Slot-based Time-Space Transformer
In the AS architecture, the temporal attention involves calculating the attention between $s_{(i,t)}$ and all previous slots. The AS structure is designed to achieve improved slot-based prediction and reasoning in complex, unguided scenarios. In previous time steps, objects were not effectively bound to specific slots, requiring each slot to search through memory to link relevant object information. The AS architecture assumes that objects were not segmented in previous frames or that effective hints for segmentation were absent.
In summary, if the upstream task effectively segments objects into slots, the CS architecture is preferred. Otherwise, the AS architecture can be considered.
For slot-based models with hints in the first frame, the AS enhancement might not be significantly effective and could increase computational load. We adopt the CS attention architecture with the T+S spatiotemporal fusion approach for our proposed STATM predictor. The memory buffer stores the slot information from the corrector for time steps. We then explain the calculation of spatiotemporal attention:
$M_{t}=\texttt{Queue}\left(S_{0},\dots,S_{t}\right).$ (2)
For a STAT encoding block, query/key/value vectors are computed for each slot:
$k_{(i,t)}^{a}$ $=LN_{k}^{a}\left(s_{(i,t)}\right)\in R^{D_{h}},$ (3a)
$q_{(i,t)}^{a}$ $=LN_{q}^{a}\left(s_{(i,t)}\right)\in R^{D_{h}},$ (3b)
$v_{(i,t)}^{a}$ $=LN_{v}^{a}\left(s_{(i,t)}\right)\in R^{D_{h}},$ (3c)
where $k$, $q$, and $v$ represent learned linear projections. $s_{(i,t)}$ denotes the vector of the $i$-th slot at time $t$. The latent dimensionality for each attention head is set to $D_{h}=D/A$.
The computation of spatiotemporal attention is also slot-based, and weights are calculated using dot-product. For the slot $s_{(i,t)}$, the spatiotemporal attention weights are computed as follows:
$a_{(i,t)}^{(a)time}$ $=\texttt{Softmax}\left(\left(\frac{q_{(i,t)}^{(a)}}{\sqrt{D_{h}}}\right)^{\mathsf{T}}\cdot\left[k_{(i,0)}^{(a)}\left\{k_{(i,t^{\prime})}^{(a)}\right\}_{t^{\prime}=1,\dots,T}\right]\right),$ (4)
$a_{(i,t)}^{(a)space}$ $=\texttt{Softmax}\left(\left(\frac{q_{(i,t)}^{(a)}}{\sqrt{D_{h}}}\right)^{\mathsf{T}}\cdot\left[k_{(0,t)}^{(a)}\left\{k_{(i^{\prime},t)}^{(a)}\right\}_{i^{\prime}=1,\dots,N}\right]\right).$
For each slot at time $t$, we calculate the weighted sum of value vectors using spatiotemporal attention coefficients from each attention head. The individual temporal or spatial attention systems in the CS structure refer to Section 3.2.
$z_{(i,t)}^{(a)time}$ $=\sum_{t^{\prime}=0}^{T}{a_{(i,t),(t^{\prime})}^{(a)time}v_{(i,t^{\prime})}^{(a)}},$ (5a)
$z_{(i,t)}^{(a)space}$ $=\sum_{i^{\prime}=0}^{N}{a_{(i,t),(i^{\prime})}^{(a)space}v_{(i^{\prime},t)}^{(a)}}.$ (5b)
The combined spatiotemporal vectors are individually linearly transformed, summed, and input into an MLP, where layer normalization (LN) is applied after each residual structure, v.i.z.,
$s_{(i,t)}^{{}^{\prime}time}$ $=W_{o}^{time}\begin{bmatrix}z_{(i,t)}^{(1)time}\\ \vdots\\ z_{(i,t)}^{(A)time}\end{bmatrix}+s_{(i,t)},$ (6d)
$s_{(i,t)}^{{}^{\prime}space}$ $=W_{o}^{space}\begin{bmatrix}z_{(i,t)}^{(1)space}\\ \vdots\\ z_{(i,t)}^{(A)space}\end{bmatrix}+s_{(i,t)},$ (6h)
$s^{\prime}_{(i,t)}$ $=\texttt{LN}\left(s_{\left(i,t\right)}^{{}^{\prime}time}+s_{\left(i,t\right)}^{{}^{\prime}space}\right),$ (7a)
$\widehat{s}_{(i,t+1)}$ $=\texttt{LN}\left(\texttt{MLP}\big{(}s^{\prime}_{(i,t)}\big{)}+s^{\prime}_{(i,t)}\right).$ (7b)
In this section, we focus on the computation process of the CS architecture using the T+S fusion approach for the STAT encoding block. In summary, temporal attention is calculated by jointly incorporating historical information from the memory buffer and spatial attention’s slot $s_{(i,t)}$. Across all structural approaches, computations are based on slots, and the equation formulations remain consistent.

### v2

@@
Abstract.
Object-centric learning aims to break down complex visual scenes into more manageable object representations, enhancing the understanding and reasoning abilities of machine learning systems toward the physical world. Recently, slot-based video models have demonstrated remarkable proficiency in segmenting and tracking objects, but they overlook the importance of the effective reasoning module. In the real world, reasoning and predictive abilities play a crucial role in human perception and object tracking; in particular, these abilities are closely related to human intuitive physics. Inspired by this, we designed a novel reasoning module called the Slot-based Time-Space Transformer with Memory buffer (STATM) to enhance the model’s perception ability in complex scenes. The memory buffer primarily serves as storage for slot information from upstream modules, the Slot-based Time-Space Transformer makes predictions through slot-based spatiotemporal attention computations and fusion. Our experimental results on various datasets indicate that the STATM module can significantly enhance the capabilities of multiple state-of-the-art object-centric learning models for video. Moreover, as a predictive model, the STATM module also performs well in downstream prediction and Visual Question Answering (VQA) tasks. We will release our codes and data at https://github.com/intell-sci-comput/STATM.
Object-Centric Learning, Slot-based Spatiotemporal Attention, Intuitive Physics, Spatiotemporal Prediction
journalyear: 2025copyright: licensedothergovconference: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.1; August 3–7, 2025; Toronto, ON, Canadabooktitle: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.1 (KDD ’25), August 3–7, 2025, Toronto, ON, Canadadoi: 10.1145/3690624.3709168isbn: 979-8-4007-1245-6/25/08ccs: Computing methodologies Computer visionccs: Computing methodologies Artificial intelligenceccs: Computing methodologies Machine learning
1. Introduction
@@
Figure 1. Slot-based Time-Space Transformer with Memory buffer architecture overview. The model employs Slot Attention (Locatello et al., 2020) for perception, which utilizes slot information predicted by STATM predictor from previous timestep and features extracted by encoder to update slot information. For the first frame, the initial slot information is obtained through either Gaussian distribution or hints module. The updated slot information is then stored in a memory buffer for subsequent use by the TATM. TATM performs reasoning by incorporating temporal cross-attention and spatial self-attention. The integration of temporal and spatial attention can be achieved in various ways. STATM supports both single-step predictions and long-sequence rollouts, where single-step prediction results can be used by Slot Attention to update slot information, and long-sequence rollout results can be used to downstream tasks such as VQA. Both perceptual and predicted slot information can be used by the decoder to obtain reconstruction results and segmentation masks. The architecture features perception and prediction modules that mutually enhance each other.
@@
The two main steps of such a model operate in a positive feedback loop. The more accurate the predictions, the better the corrections become. Consequently, the more accurate the corrections, the more precise the information provided for the prediction step is, leading to better predictions. Therefore, having a reasonable and efficient predictor is crucial for the entire model. In real-world scenarios, humans also engage in prediction as a crucial aspect of their object perception and tracking, but their prediction behaviors often involve more intricate processes. Humans typically combine the motion state of an object with the interactions of other objects to predict possible future states and positions of the object. The object’s motion state is inferred by humans using their common sense from the object’s past positions over a while. In so doing, humans enhance their ability to recognize and track relevant objects within complex scenes, which is an integral component of human intuitive physics (Sudderth, 2006; Ullman et al., 2017; Mitko and Fischer, 2020). The prediction step in SAVi and SAVi++ is similar to human inference, but their predictor module is somewhat simplistic, as it relies solely on single-frame information from the current time step for prediction.
In the field of object-centric video prediction, SlotFormer (Wu et al., 2023) and OCVP (Villar-Corrales et al., 2023) transform the spatiotemporal attention structure used for video classification in TimeSformer (Bertasius et al., 2021) into a similar structure utilized for slot prediction. However, the number of slots (also considered as tokens) required by SlotFormer is likely to increase dramatically with time as the number of objects in the scene grows, leading to an excess of superfluous slot computation. OCVP explored two types of spatiotemporal slot models, which, to some extent, mitigated the increase in token numbers, yet still faced the issue of unnecessary slot computations. The predictive capability of both approaches heavily depends on the quality of the upstream slot extraction. Neither approach made improvements to the upstream module responsible for slot extraction, nor did they delve into the impact of prediction on upstream perception.
@@
Human perception and prediction are typically complementary. When assessing an object’s movement, humans often rely on short-term memory impressions of the object. These impressions, along with consideration of environmental factors within the scene, are used to predict the object’s movement. Through a comprehensive analysis of time and space, humans anticipate the object’s next position. Drawing inspiration from human behavior, we introduce a novel prediction module aimed at enhancing slot-based models for video. This module comprises two key components: 1) Slot-based Memory Buffer: designed to store historical slot information obtained from the upstream module. 2) Slot-based Time-Space Transformer Module: designed by applying spatiotemporal attention mechanisms to slots for inferring the temporal motion states of objects and calculating spatial objects interactions, which also integrates time and space attention results. The module only computes using slots from the current moment and those from past moments. This not only addresses the issue of the increasing number of tokens as time progresses and the number of objects in the scene increases, but it also reduces unnecessary slot computations. We term the proposed model as Slot-based Time-Space Transformer with Memory buffer (STATM). Upon substituting the prediction module of SAVi and SAVi++ into the STATM, we observe distinct impacts of different spatiotemporal fusion methods on SAVi and SAVi++. By employing an appropriate fusion method and memory buffer sizes, we observed a significant enhancement in the object segmentation and tracking capabilities of SAVi and SAVi++ on videos containing complex backgrounds and multiple objects per scene.
Overall, our contributions are summarized as follows:
  • We have investigated the impact of prediction modules on models utilizing Slot Attention (Locatello et al., 2020) for object-centric learning from video and have developed the STATM module as a predictor. By simply incorporating a memory buffer and spatiotemporal attention, we have significantly enhanced the capabilities of models like SAVi (Kipf et al., 2021), and SAVi++(Elsayed et al., 2022).
  • We have diligently worked to reduce the computational cost of the spatiotemporal module. In contrast to other models that utilize multiple frames for prediction (Wu et al., 2023; Villar-Corrales et al., 2023), our spatiotemporal module only combines current and past slots for computation, effectively decreasing the number of tokens required for the prediction module and the amount of slot computations it performs. As time progresses and the number of objects in the scene increases, the advantage of STATM becomes even more apparent.
  • We have conducted experiments across multiple benchmarks. We have observed that the STATM module significantly enhances the capabilities of multiple state-of-the-art object-centric learning models for video, such as SAVi (Kipf et al., 2021) , SAVi++ (Elsayed et al., 2022), and STEVE (Singh et al., 2022). Moreover, as a predictive model, the STATM also performs well in downstream prediction and Visual Question Answering (VQA) tasks.
  • Furthermore, we have briefly explored the impact of various spatiotemporal architectures and different memory buffer sizes on model performance.
2. Related Work
@@
Object-centric Learning. In recent years, object-centric learning has emerged as a significant research direction in computer vision and machine learning. It aims to enable machines to perceive and understand the environment from an object-centered perspective, thereby constructing more intelligent visual systems. There is a rich literature on this research, including SQAIR (Kosiorek et al., 2018), R-SQAIR (Stanić and Schmidhuber, 2019), SCALOR (Jiang et al., 2019), Monet (Burgess et al., 2019), OP3 (Veerapaneni et al., 2020), ViMON (Weis et al., 2020), PSGNet (Bear et al., 2020), SIMONe (Kabra et al., 2021), and others (Kahneman et al., 1992; Kipf et al., 2019; Zhang et al., 2022; Xie et al., 2022). Slot-based Models represent a prominent approach within object-centric learning. They achieve this by representing each object in a scene as an individual slot, which is used to store object features and attributes (Kumar et al., 2020; Yang et al., 2021; Dittadi et al., 2021; Hassanin et al., 2022; Wu et al., 2023).
@@
Slot-based Attention and spatiotemporal Attention. Our current work is closely related to slot-based attention and spatiotemporal attention. There are a lot of works related to slot-based attention (Locatello et al., 2020; Hu et al., 2020; Zoran et al., 2021; Ye et al., 2021; Wang et al., 2023; Wu et al., 2023). Spatiotemporal attention mechanisms are particularly effective in handling video data or time-series data, allowing networks to understand and leverage relationships between different time steps or spatial positions (Li et al., 2020; Bertasius et al., 2021; Luo et al., 2021). Currently, they find wide applications in various fields such as video object detection and tracking (Lin et al., 2021; Chen et al., 2022), action recognition (Yang et al., 2022), natural language processing (Xu et al., 2020; Weld et al., 2022), medical image processing (Zhang et al., 2020), among many others (Ding et al., 2020; Yuan et al., 2020; Cheng et al., 2020).
Prediction and Inference on Physics. The implementation of object-centric physical reasoning is crucial for human intelligence and is also a key objective in artificial intelligence. Interaction Network (Battaglia et al., 2016) as the first general-purpose learnable physics engine, is capable of performing reasoning tasks centered around objects or relationships. Another similar study is the Neural Physics Engine (Chang et al., 2016). On the other hand, Visual Interaction Networks (Watters et al., 2017) can learn physical laws from videos to predict the future states of objects. Additionally, there are many models developed based on this foundation (Chen et al., 2021a; Jusup et al., 2022; Meng et al., 2022; Piloto et al., 2022; Singh et al., 2022; Driess et al., 2023). Additionally, there are many object-centric predictive models that are based on slot representations (Ding et al., 2021a, b; Wu et al., 2023). However, their performance largely depends on the quality of slot extraction by upstream perception modules. In order to achieve a deeper understanding of commonsense intuitive physics within artificial intelligence systems, (Piloto et al., 2022) have built a system capable of learning various physical concepts, albeit requiring access to privileged information such as segmentation. Our research primarily aims to construct an object-centric system for object perception, learning of physics, and reasoning.
@@
3. Slot-based Time-Space Transformer with Memory Buffer (STATM)
@@
To enhance the slot-based video models, we introduce a new module called Slot-based Time-Space Transformer with Memory Buffer (STATM) as the predictor. This module consists of two key components: 1) memory buffer, and 2) Slot-based Time-Space Transformer (STAT). The memory buffer serves as a repository for storing historical slot information obtained from upstream modules, while STAT utilizes the information stored in the memory buffer for prediction and causal reasoning. The overall framework is shown in Figure 1.
3.1. Memory Buffer
@@
3.2. Slot-based Time-Space Transformer
The primary role of STAT (Slot-based Time-Space Transformer) lies in leveraging slot data from the memory buffer to facilitate slot-based dynamic temporal reasoning and spatial interaction computations. Furthermore, it integrates the outcomes of temporal reasoning and spatial interactions to achieve a unified understanding. Specifically, for temporal dynamic reasoning, a cross-attention mechanism is employed, which effectively utilizes historical context stored in the memory buffer to enable accurate predictions of future states. Meanwhile, for spatial interaction computations, we employ a self-attention mechanism that operates on slot representations to compute the relevance between different slots within the $S$. The results obtained from temporal dynamic reasoning and spatial interaction computation are merged to provide a holistic understanding encompassing both temporal dynamics and spatial interactions. This comprehensive representation enhances the model’s capability for accurate prediction and reasoning in object-centric tasks. We propose three approaches:
(2a) $\widehat{S}_{t+1}$ $=CrossAtt^{time}(S_{t},M_{t})+SelfAtt^{space}(S_{t})$
(2b) $\widehat{S}_{t+1}$ $=CrossAtt^{time}(SelfAtt^{space}(S_{t}),M_{t})$
(2c) $\widehat{S}_{t+1}$ $=SelfAtt^{space}(CrossAtt^{time}(S_{t},M_{t})),$
(2a) T+S: The sum of computed temporal attention and spatial attention. (2b) ST: Spatial attention computation followed by using the outcome as input for temporal attention. (2c) TS: Temporal attention computation followed by using the outcome as input for spatial attention.
Figure 2. Spatiotemporal attention computation architectures. The green slots represent those employed for spatial attention computation, while the orange slots are indicative of those used for temporal attention computation.
As shown in Figure 2, we introduce two computational architectures for spatiotemporal attention: (a) Corresponding Slot Attention (CS): For slot $s_{(i,t)}$, temporal attention is computed by using it and corresponding slots in $\{s_{(i,0)},\dots,s_{(i,t-1)}\}$, while spatial attention computation is performed using it and all slots within $\{s_{(0,t)},\dots,s_{(N,t)}\}$. (b) All Slot Attention (AS): For slot $s_{(i,t)}$, temporal attention is computed by using it and all slots in $\{s_{(0,0)},\dots,s_{(N,t-1)}\}$. The spatial attention computation remains the same as in the CS.
@@ 3.2. Slot-based Time-Space Transformer
In the AS architecture, the temporal attention involves calculating the attention between $s_{(i,t)}$ and all previous slots.
3.3. Differences between STATM and Other Slot-based Models
In the field of object-centric video prediction, SlotFormer (Wu et al., 2023) and OCVP (Villar-Corrales et al., 2023) transform the spatiotemporal attention structure used for video classification in TimeSformer (Bertasius et al., 2021) into a similar structure utilized for slot prediction. Nevertheless, the computational slots utilized by these models are likely to surge dramatically over time as the number of objects in the scene increases, leading to an excess of unnecessary slot computations.
Our model design is focused on reducing the computational cost. Traditionally, slot-based models, such as SlotFormer(Wu et al., 2023), flatten the time $T$ ($T>2$) and the number of slots $N$ ($N>2$) when using $T$ time steps for prediction, then input them into the transformer encoder to calculate full attention between each slot. Thus, the transformer encoder must process $T\times N$ tokens and perform attention calculations on $T\times N$ slots, leading to $(T\times N)^{2}$ calculations.
In contrast, our spatiotemporal modules only combine the current time slots with the previous time slots for calculation. Therefore, the prediction modules based on the CS structure only requires $T+N$ tokens, and the AS structure requires $T\times N$ tokens. In attention calculations, we only consider computations between the current moment slots and other slots. For the CS structure, in terms of time attention, it only needs to calculate attention between current time slot $s_{i}$ ($i=1,\dots,N$) and the other $T-1$ slots, and in terms of space attention, it only needs to calculate attention between $N$ slots, thus involving $N\times(T-1)+N^{2}$ slot calculations in total. For the AS structure, it involves $(T-1)\times N\times N$ calculations in time and $N^{2}$ calculations in space, totaling $T\times N^{2}$ calculations. Clearly, aside from the token requirement for the AS structure, both the number of tokens and the amount of computation required by our spatiotemporal attention module are significantly less than that of SlotFormer. We also found that using only the CS structure can achieve good model performance. As the number of frames and objects in a scene increase, the CS structure is more efficient.
3.4. Training
Our focus is on conducting perception and prediction experiments. In the perception experiments, we intend to enhance several state-of-the-art object-centric learning models using the STATM module. Consequently, the training loss for each model may differ. We train STATM-SAVi and STATM-SAVi++ to minimize the L2 loss between the predicted and ground-truth targets, such as optical flow, images, and depth signals. The training loss for STATM-STEVE aligns with the loss function used in STEVE(Singh et al., 2022).
In the prediction experiments, we train the STATM model by jointly minimizing slots and images reconstruction loss (also L2).
4. Experiments
Our experiments primarily comprise three parts: perception, prediction and VQA, and ablation. The perception experiments aim to verify the impact of the STATM module on existing state-of-the-art object-centric learning models for video. The prediction experiments are designed to preliminarily demonstrate the robust performance of the STATM module as a predictive model in downstream prediction and VQA tasks. The ablation studies focus on assessing the effects of the memory buffer and various spatiotemporal structures on model performance.
Baselines. In the perception experiments, we primarily compare SAVi (Kipf et al., 2021), SAVi++ (Elsayed et al., 2022), STEVE (Singh et al., 2022) and SAVi-SlotFormer. For SAVi, we chose official implementation, SAVi-small, as the baseline. The SAVi-samll model includes five components: encoder, decoder, slot initialization, corrector, and predictor. The encoder uses a CNN to extract features from video frames. Slot initialization, using either an MLP or a CNN, prepares slots with initial data like bounding boxes. The corrector, powered by Slot Attention (Locatello et al., 2020), updates slots using encoder features. The predictor, a transformer block, uses self-attention for forecasting and initializes the corrector for consistent tracking. Finally, the decoder outputs RGB predictions and an alpha mask using a Spatial Broadcast Decoder. The SAVi-SlotFormer is a baseline we develope to assess the impact of the prediction module on the perceptual performance of SAVi. SAVi++ (Elsayed et al., 2022) has a structure similar to SAVi with a ResNet34 backbone. Unlike SAVi, SAVi++ introduces depth as self-supervised objectives. It also incorporates data augmentation and utilizes a transformer encoder after ResNet34.
In the prediction section, we compare G-SWM (Lin et al., 2020), SAVi-dyn(Wu et al., 2023), and SlotFormer (Wu et al., 2023). SlotFormer (Wu et al., 2023) is a transformer-based framework for object-centric visual simulation. It leverages slots extracted by upstream modules like SAVi to train a slot-based transformer encoder model for prediction purposes.
For the VQA experiments, our main comparisons involve DCL (Chen et al., 2021b), VRDP (Ding et al., 2021a), and SlotFormer (Wu et al., 2023). SlotFormer utilizes prediction results from rollout simulations to train Aloe (Ding et al., 2021b) for VQA tasks. VRDP (Ding et al., 2021a) is designed to jointly learn visual concepts and infer physics models of objects and their interactions from both videos and language. It primarily consists of three modules: a visual perception module, a concept learner, and a differentiable physics engine. We have implemented VRDP with a visual perception module that is trained based on object properties.
For more baselines, please refer to Appendix Section  A.
Metrics. To evaluate the model’s object-centric learning capability for video, we selected the Adjusted Rand Index (ARI) (Rand, 1971; Hubert and Arabie, 1985) and the mean Intersection over Union (mIoU) as evaluation metrics. ARI quantifies the alignment between predicted and ground-truth segmentation masks. For scene decomposition assessment, we commonly employ FG-ARI and, which is a permutation-invariant clustering similarity metric. It allows us to compare inferred segmentation masks to ground-truth masks while excluding background pixels. mIoU is a widely used segmentation metric that calculates the mean Intersection over Union values for different classes or objects in a segmentation task. To assess video quality, we report PSNR, SSIM (Wang et al., 2004), and LPIPS (Zhang et al., 2018). To evaluate the prediction outcomes, we utilize FG-ARI and FG-mIoU.
Datasets. To evaluate the object-centric learning capability, we utilized the synthetic Multi-Object Video (MOVi) datasets (Research, 2020; Greff et al., 2022). These datasets are divided into five distinct categories: A, B, C, D, and E. MOVi-A and B depict relatively straightforward scenes, each containing a maximum of 10 objects. MOVi-C, D, and E present more intricate scenarios with complex natural backgrounds. MOVi-C, generated using a stationary camera, presents scenes with up to 10 objects. Transitioning to MOVi-D, the dataset extends the object count to accommodate a maximum of 23 objects. Lastly, MOVi-E introduces an additional layer of complexity by incorporating random linear camera movements. Each video sequence is sampled at a frame rate of 12, resulting in a total of 24 frames per second.
To assess the predictive and Visual Question Answering (VQA) capabilities, we have selected the CLEVRER (Yi et al., 2019) dataset. CLEVRER dataset is specifically designed for video understanding and reasoning, focusing on the dynamics of objects and their causal interactions. For the VQA task, CLEVRER incorporates four types of questions: descriptive (e.g., ”what color”), explanatory (”what’s responsible for”), predictive (”what will happen next”), and counterfactual (”what if”). The predictive questions require the model to simulate future object interactions, such as collisions. Thus, we are particularly concentrating on enhancing the accuracy of predictive questions through the implementation of STATM’s future rollout.
Training Setup. In all experiments except the ablation study in Section 4.3, we used the STAT encoding block in combination with the CS attention architecture, featuring the T+S spatiotemporal fusion approach. For perception experiments, we utilized videos comprising of 6 frames at a resolution of 64$\times$64 pixels to train the STATM-SAVi and SAVi models. The training process is conducted over 100k iterations. Similarly, the STATM-SAVi++ and SAVi++ models were trained on continuous videos consisting of 6 frames at a higher resolution of 128$\times$128 pixels, with training duration encompassing 100k or 500k iterations. The buffer size was unconstrained during training, and the maximum length of effective information was limited to 6 due to the utilization of a 6-frame training sequence. Bounding boxes were used as the conditioning for all models. For prediction and VQA experiments, we train our models (STATM-SAVi) for 400k steps with a batch size of 64 on the CLEVRER dataset to extract slots. The number of slots is set to 7, with a learning rate of 0.0001. We subsample the video by a factor of 2 to train STATM, conducting approximately 500k training steps with a batch size of 64 and a learning rate of 0.0002. We use rollout slots to train Aloe (Ding et al., 2021b), targeting around 300k steps with a learning rate of 0.0001 and a batch size of 128. We use the Adam optimizer and apply warm-up and decay learning rate schedule for the first 2.5% of the total training steps. For more training setup, please refer to Appendix Section  A and  B.
4.1. Perception